# Towards exponentially-convergent simulations of extreme-mass-ratio inspirals: A time-domain solver for the scalar Teukolsky equation with singular source terms

### v2

@@ IV.2.1 Price tails on Schwarzschild
and where $\phi_{\ell m}=\partial_{r_{*}}\psi_{\ell m}$. We solve Eq. (II.4) (with $\ell=2$ and $M=1$) on $\rho\in[-200,1200]$ with the layer at $R=150$. We use $K=400$ subdomains and on each subdomain the solution is approximated by a degree $N=10$ polynomial. The time step is set to $\Delta t=0.0866$. In Fig. 3, we show the solution recorded as a time series at $r_{*}=500$ (left panel) and at $\mathscr{I}^{+}$ (right panel). At late times we empirically measure the decay rate of solution and quote this value in the parentheses of the figure’s label. In the inset figure, we also report the local power index and show the expected rate (as a horizontal line) based on Price’s Law Price (1972). Our measured tail decay rates are consistent with the theoretically known values at both a finite distance and $\mathscr{I}^{+}$ Hod (1999).
@@ IV.2.2 Price tails on Kerr
Table 1: We reproduce the $(\ell^{\prime},m)\rightarrow(\ell,m)$ scattering matrix from Ref. Burko and Khanna (2011) where they verified the late time tail behavior of scalar fields in the Kerr background using quad precision. We set the spin on the primary black hole as $a=0.8$. The expected behavior is $\ell^{\prime}+\ell+1$ if $\ell<\ell^{\prime}$ and $\ell^{\prime}+\ell+3$ if $\ell\geq\ell^{\prime}$. We get the expected behavior with less than $1\%$ of error. Since the higher modes are very weakly coupled to the initial $\ell^{\prime}$ mode, double-precision round-off errors prevent us from obtaining accurate computations of the power-law behavior for certain couplings. More accurate results can be obtained with quad-precision floating-point computations, which we hope to consider in the future.
@@ IV.3.1 Fluxes on Schwarzschild
We first consider the simpler Schwarzschild case, where $a=0$ and the harmonic modes decouple for all values of $(\ell,m)$. For our simulations, we have typically chosen $M=1$, and we solve Eq. (II.4) with a distributional source term given by Eq. (14). Our physical domain $\rho\in[-100,200]$ is partitioned into 200 subdomains and on each subdomain we approximate the solution with a degree $N=10$ polynomial. We set the final time $T=8000$, and the time step $\Delta t=0.0018$. As is common practice, we supply trivial (zero) initial data. We note that supplying trivial initial data is commonly done because (i) the exact initial data is unknown and (ii) the expectation is that over long enough times the impact of incorrect initial data will propagate away as so-called junk radiation Field et al. (2010). Trivial initial data is inconsistent with the distributional forcing terms (14), so we smoothly turn on the source term with a ramp-up function. We take our ramp-up function to be Eq. C1 of Ref. Field et al. (2009) with parameter values $\tau=400$ and $\delta=0.00025$. To achieve high accuracies for some of the higher harmonic modes, we sometimes deviate from these default settings.
@@ V Summary and future work
As a byproduct of our work, we have compiled numerical evidence that when combined with the hyperboloidal layer method, one seemingly reasonable choice of the first-order reduction variable appears to be problematic. This issue shows up in different settings, such as calculating scalar fluxes at future null infinity or when comparing to exact solutions of the ordinary wave equation. In all settings, the numerical solution’s accuracy saturates at about 5 to 6 digits irrespective of the timestep or grid resolution. On the other hand, with a better choice of auxiliary variable, we are able to achieve the expected decay of numerical error to better than 12 digits of accuracy.
@@ Acknowledgments
We thank Anil Zenginoglu for discussions on first-order reductions and hyperboloidal layers, Jennifer Ryan for discussions on superconvergence, and Olivier Sarbach for helpful feedback on App. C. We thank Som Bishoyi, Peter Diener, Zachariah Etienne, Alfa Heryudono, Scott A. Hughes, Harald Pfeiffer, Leo Stein, Saul Teukolsky, Alex Vano-Vinuales, Niels Warburton, and Barry Wardell for helpful discussions throughout the project. We also thank Karoly Csukas and Hannes Rüter for their valuable feedback on an earlier draft of this paper. The authors acknowledge support of NSF Grants PHY-2106755 and PHY-2307236 (G.K), DMS-1912716 and DMS-2309609 (S.F, S.G, and G.K), AFOSR Grant No. FA9550-18-1-0383 (S.G), PHY-2207780 (K.R), and Office of Naval Research/Defense University Research Instrumentation Program (ONR/DURIP) Grant No. N00014181255. MV acknowledges partial support through the cyberteams CAREER’s program, NSF Grant No. OAC-2018873. K.R is a member of the Weinberg Institute and this manuscript has preprint number WI-45-2023. This material is based upon work supported by the National Science Foundation under Grant No. DMS-1439786 while a subset of the authors were in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, RI, during the Advances in Computational Relativity program. Some of the simulations were performed on the UMass-URI UNITY supercomputer supported by the Massachusetts Green High Performance Computing Center (MGHPCC). We also thank ChatGPT for writing the first draft of the conclusion (and only the conclusion) of this paper.
@@ Appendix A Mode-coupling coefficients
When expanding the $s=0$ Teukolsky equation (1) in spherical harmonics, we encounter a term $\sin^{2}\theta Y_{\ell m}$ that is responsible for mode coupling. To re-expand this term with spherical harmonics
@@
Appendix B Matrix form of the full system (II.4)
We now consider the matrix form for the general system (II.4) for any combination of $\ell_{\rm max}$, $m$ and $\ell$-mode type = {odd, even}. The even and odd $\ell$ modes decouple so their governing systems can be analyzed separately. For concreteness, consider (for a fixed value of $m$) the even $\ell$-mode sector, where $\ell=\ell_{\rm min},\ell_{\rm min}+2,\dots,\ell_{\rm max}$. First, we split our solution vector $U_{\ell_{\rm max}m}$ into subvectors
@@
where $E_{\pi\pi}^{-1}E_{\pi\pi}=\mathbf{I}$. We see that in general $A$ is not a symmetric matrix, even though $E_{\pi\pi}^{-1}$ is. By considering $H=0$ we see that each of the four blocks of $A$ are symmetric when working in the original $t$ and $r_{*}$ coordinates. As a special case, $A$ is symmetric when $a=0$.
Appendix C The first-order Eq. (II.4) is a positive (semi-definite) symmetric hyperbolic system
With an explicit expression for the matricies of our problem given in App. B, we now show conditions under which the first-order system (II.4) is symmetric hyperbolic. Recall that to study a first-order system’s hyperbolicity classification it is sufficient to consider the principle part of the differential equations,
@@
where the variable-coefficient matrices $A(\rho)$, $E(\rho)$, and $\hat{A(\rho)}$ depend on the spatial independent variable $\rho$ and $\dots$” denotes lower-order terms (ie. it does not include any term proportional to $\dot{U}$ or $U^{\prime}$). Eq. (75) represents the form of our original system (II.4)’s principle part after we put it into matrix form (25) where the matrices are given in App. B. System (76) is the result of inverting $E$, leading to the system (26) that we discretize with the DG method.
@@
where $U^{T}$ is the transpose of $U$. Then ${\cal E}(\tau)\geq 0$ due to $E(\rho)$ being positive (semi-)definite. Our approach is as follows: provided the equality between norms of $U$ and $V$ in Eq. (78) hold, the problem for $U$ is well-posed if we show the problem for $V$ is well-posed. Well-posedness for $V$ follows from it being symmetric hyperbolic: we show the system for $U$ is positive symmetric at all points in the physical domain except at the right-most boundary point $\rho=s$ (from which it follows the system for $V$ is symmetric hyperbolic except at $\rho=s$), and at $\rho=s$ the system for $V$ is symmetric hyperbolic by direct inspection.
@@ C.1 Setup
The system of first-order differential equations (II.4) are solved on an open spatial domain, where $0\leq a\leq 1$ and $a=1$ corresponds to an extremal Kerr blackhole. In the original $(t,r_{*})$ coordinate system ($\Omega=1$ and $H=0$) this is $r_{*}\in[\rho_{L},\infty)$ and we note that $r_{*}=-\infty$ is the black hole’s horizon. In hyperboloidal coordinates $(\tau,\rho)$ the domain is $\rho\in[\rho_{L},s]$ where $r_{*}(\rho=s)=\infty$.
@@ C.2 $a=0$ with hyperboloidal layers
We first consider a special case: when $a=0$ (scalar waves on Schwarzschild) the matrix $E$ is positive definite for all $\rho\in D\setminus\{s\}$. Upon further setting $M=0$, the results of this subsection also apply to the ordinary flatspace wave equation in hyperbolodial coordinates (IV.1).
@@ C.3 The general case
We first show the diagonal entries are positive, which in turn require that the diagonal entries of both $E_{\phi\phi}$ and $E_{\pi\pi}$ are positive. Because $E_{\phi\phi}=\mathbf{I}$ this is obviously true. Next we consider the diagonal of $E_{\pi\pi}$, whose elements are of the form $1-fc^{\ell}_{\ell m}$ where $f(r;a)$ is defined in Eq. (17). Let us first remark that $\Delta>0$ (and hence $f>0$) outside of the horizon, which our computational grid is restricted to. From App. A recall that $10<c^{\ell}_{\ell m}<$ 1, and so $1-f<1-fc^{\ell}_{\ell m}$. At this point, we use Mathematica to find an upper bound on $f(r;a)$. We do this by first finding $r=r_{\mathrm{max}}$ that maximizes $f(r;a)$ for a fixed value of $a$. Setting $\partial_{r}f=0$, we find $r_{\mathrm{max}}$ solves the cubic equation $r^{3}-3Mr^{2}+a^{2}r+a^{2}M=0$, which has one real root outside of the horizon. We now view $f(r=r_{\mathrm{max}};a)$ as a function of $a$. Numerical evidence (and physical intuition) suggests that this function is maximized at $a=1$. In fact, for $a=M=1$ we have $r_{\mathrm{max}}=1+\sqrt{2}$. Evaluating we get
@@ C.4 System at infinity
And so the norms of $C^{-1}\hat{A}C^{-1}$ and $A$ are identical at every point $\rho\in D$. In particular, at $\rho=s$ we have that $C^{-1}\hat{A}C^{-1}$ is bounded because $A$ is bounded. Therefore $C^{-1}\hat{A}C^{-1}$ exists, and hence the associated system (77) is symmetric hyperbolic everywhere in the the physical domain of interest including $\rho=s$.
@@ Appendix D An alternative first-order reduction
Throughout the main body of this paper, we have worked with a first-order reduction of Eq. (II.1) that arise from the reduction variables $\pi_{\ell m}=-\frac{\partial\psi_{\ell m}}{\partial t}$ and $\phi_{\ell m}=\frac{\partial\psi_{\ell m}}{\partial r_{*}}$ followed by a coordination transformation 18. Appendix C shows the conditions under which the resulting system is positive (semi-definite) symmetric hyperbolic.
@@
where in carrying out the first-order reduction, we replaced the mixed-partial derivative term as $2\dot{\psi}^{\prime}_{\ell m}=-\partial_{\rho}\pi_{\ell m}+\partial_{\tau}\phi_{\ell m}$ and did not consider alternative choices. For example, the choice $2\dot{\psi}^{\prime}_{\ell m}=-2\partial_{\rho}\pi_{\ell m}$ will lead to a simpler $A$ matrix (cf. 25) but was not considered in our first-order reduction. If $\psi_{\ell m}$ solves the first-order system (D) it also solves the original second-order equation (D) provided $\phi_{\ell m}=\partial_{\rho}\psi_{\ell m}$.
Let us now put Eq. (D) into matrix form (25). For concreteness, consider the even $\ell$-mode sector with $m=0$ and $\ell_{\rm max}=2$. Then we have $U=\left[\pi_{00},\pi_{20},\phi_{00},\phi_{20}\right]^{T}$, and
@@
Numerically, we have been unable to achieve good results with this system. For example, in the very simple case of $M=0$ (ie flatspace) the numerical solution to this system stops converging after about 4 to 5 digits of accuracy. This is shown most clearly in Fig. 2: compare “System 2” (the problematic one using reduction variables (86)) and “System 1” (the first-order system that arise from the reduction variables (15)).

### v3

@@ IV.2.1 Price tails on Schwarzschild
and where $\phi_{\ell m}=\partial_{r_{*}}\psi_{\ell m}$. We solve Eq. (20) (with $\ell=2$ and $M=1$) on $\rho\in[-200,1200]$ with the layer at $R=150$. We use $K=400$ subdomains and on each subdomain the solution is approximated by a degree $N=10$ polynomial. The time step is set to $\Delta t=0.0866$. In Fig. 3, we show the solution recorded as a time series at $r_{*}=500$ (left panel) and at $\mathscr{I}^{+}$ (right panel). At late times we empirically measure the decay rate of solution and quote this value in the parentheses of the figure’s label. In the inset figure, we also report the local power index and show the expected rate (as a horizontal line) based on Price’s Law Price (1972). Our measured tail decay rates are consistent with the theoretically known values at both a finite distance and $\mathscr{I}^{+}$ Hod (1999).
@@ IV.2.2 Price tails on Kerr
Table 1: We reproduce the $(\ell^{\prime},m)\rightarrow(\ell,m)$ scattering matrix from Ref. Burko and Khanna (2011) where they verified the late time tail behavior of scalar fields in the Kerr background using quad precision. We set the spin on the primary black hole as $a=0.8$. The expected behavior is $\ell^{\prime}+\ell+1$ if $\ell<\ell^{\prime}$ and $\ell^{\prime}+\ell+3$ if $\ell\geq\ell^{\prime}$. We get the expected behavior with less than $1\%$ of error for the lower =$\ell=\{0,1,2,3\}$ modes. Since the higher modes are very weakly coupled to the initial $\ell^{\prime}$ mode, double-precision round-off errors prevent us from obtaining similar $1\%$ accuracy computations of the power-law behavior for some of the higher =$\ell=\{4,5\}$ modes. More accurate results can be obtained with quad-precision floating-point computations, which we hope to consider in the future.
@@ IV.3.1 Fluxes on Schwarzschild
We first consider the simpler Schwarzschild case, where $a=0$ and the harmonic modes decouple for all values of $(\ell,m)$. For our simulations, we have typically chosen $M=1$, and we solve Eq. (20) with a distributional source term given by Eq. (14). Our physical domain $\rho\in[-100,200]$ is partitioned into 200 subdomains and on each subdomain we approximate the solution with a degree $N=10$ polynomial. We set the final time $T=8000$, and the time step $\Delta t=0.0018$. As is common practice, we supply trivial (zero) initial data. We note that supplying trivial initial data is commonly done because (i) the exact initial data is unknown and (ii) the expectation is that over long enough times the impact of incorrect initial data will propagate away as so-called junk radiation Field et al. (2010). Trivial initial data is inconsistent with the distributional forcing terms (14), so we smoothly turn on the source term with a ramp-up function. We take our ramp-up function to be Eq. C1 of Ref. Field et al. (2009) with parameter values $\tau=400$ and $\delta=0.00025$. To achieve high accuracies for some of the higher harmonic modes, we sometimes deviate from these default settings.
@@ V Summary and future work
As a byproduct of our work, we have compiled numerical evidence that when combined with the hyperboloidal layer method, one seemingly reasonable choice of the first-order reduction variable appears to be problematic. This issue shows up in different settings, such as calculating scalar fluxes at future null infinity or when comparing to exact solutions of the ordinary wave equation. In all settings, the numerical solution’s accuracy saturates at about 5 to 6 digits irrespective of the timestep or grid resolution. On the other hand, with a better choice of auxiliary variable, we are able to achieve the expected decay of numerical error to better than 12 digits of accuracy. Appendices C and D discuss the hyperbolicity and well-posedness properties of both system.
@@ Acknowledgments
We thank Stephen Lau and Anil Zenginoglu for discussions on first-order reductions and hyperboloidal layers, Jennifer Ryan for discussions on superconvergence, and Olivier Sarbach for helpful feedback on App. C. We thank Som Bishoyi, Peter Diener, Zachariah Etienne, Alfa Heryudono, Scott A. Hughes, Harald Pfeiffer, Leo Stein, Saul Teukolsky, Alex Vano-Vinuales, Niels Warburton, and Barry Wardell for helpful discussions throughout the project. We also thank Karoly Csukas and Hannes Rüter for their valuable feedback on an earlier draft of this paper. The authors acknowledge support of NSF Grants PHY-2106755 and PHY-2307236 (G.K), DMS-1912716 and DMS-2309609 (S.F, S.G, and G.K), AFOSR Grant No. FA9550-18-1-0383 (S.G), PHY-2207780 (K.R), and Office of Naval Research/Defense University Research Instrumentation Program (ONR/DURIP) Grant No. N00014181255. MV acknowledges partial support through the cyberteams CAREER’s program, NSF Grant No. OAC-2018873. K.R is a member of the Weinberg Institute and this manuscript has preprint number WI-45-2023. This material is based upon work supported by the National Science Foundation under Grant No. DMS-1439786 while a subset of the authors were in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, RI, during the Advances in Computational Relativity program. Some of the simulations were performed on the UMass-URI UNITY supercomputer supported by the Massachusetts Green High Performance Computing Center (MGHPCC). We also thank ChatGPT for writing the first draft of the conclusion (and only the conclusion) of this paper.
@@ Appendix A Mode-coupling coefficients
When expanding the $s=0$ Teukolsky equation (II.1) in spherical harmonics, we encounter a term $\sin^{2}\theta Y_{\ell m}$ that is responsible for mode coupling. To re-expand this term with spherical harmonics
@@
Appendix B Matrix form of the full system (20)
We now consider the matrix form for the general system (20) for any combination of $\ell_{\rm max}$, $m$ and $\ell$-mode type = {odd, even}. The even and odd $\ell$ modes decouple so their governing systems can be analyzed separately. For concreteness, consider (for a fixed value of $m$) the even $\ell$-mode sector, where $\ell=\ell_{\rm min},\ell_{\rm min}+2,\dots,\ell_{\rm max}$. First, we split our solution vector $U_{\ell_{\rm max}m}$ into subvectors
@@ Appendix B Matrix form of the full system (20)
where $E_{\pi\pi}^{-1}E_{\pi\pi}=\mathbf{I}$. We see that, in general, $A$ is not a symmetric matrix, even though $E_{\pi\pi}^{-1}$ is. By considering $H=0$ we see that each of the four blocks of $A$ are symmetric when working in the original $t$ and $r_{*}$ coordinates. As a special case, $A$ is symmetric when $a=0$.
Appendix C The first-order Eq. (20) is a positive (semi-definite) symmetric hyperbolic system
With an explicit expression for the matricies of our problem given in App. B, we now show conditions under which the first-order system (20) is symmetric hyperbolic. Recall that to study a first-order system’s hyperbolicity classification it is sufficient to consider the principle part of the differential equations,
@@
where the variable-coefficient matrices $A(\rho)$, $E(\rho)$, and $\hat{A(\rho)}$ depend on the spatial independent variable $\rho$ and $\dots$” denotes lower-order terms (ie. it does not include any term proportional to $\dot{U}$ or $U^{\prime}$). Eq. (59) represents the form of our original system (20)’s principle part after we put it into matrix form (25) where the matrices are given in App. B. System (60) is the result of inverting $E$, leading to the system (26) that we discretize with the DG method.
@@
where $U^{T}$ is the transpose of $U$. Then ${\cal E}(\tau)\geq 0$ due to $E(\rho)$ being positive (semi-)definite. Our approach is as follows: provided the equality between norms of $U$ and $V$ in Eq. (62) hold, the problem for $U$ is well-posed if we show the problem for $V$ is well-posed. Well-posedness for $V$ follows from it being symmetric hyperbolic: we show the system for $U$ is positive symmetric at all points in the physical domain except at the right-most boundary point $\rho=s$ (from which it follows the system for $V$ is symmetric hyperbolic except at $\rho=s$), and at $\rho=s$ the system for $V$ is symmetric hyperbolic by direct inspection. Caveats about strong well-posedness, which are applicable to the first-order system (20) as well, are summarized in App. D.1.
@@ C.1 Setup
The system of first-order differential equations (20) are solved on an open spatial domain, where $0\leq a\leq 1$ and $a=1$ corresponds to an extremal Kerr blackhole. In the original $(t,r_{*})$ coordinate system ($\Omega=1$ and $H=0$) this is $r_{*}\in[\rho_{L},\infty)$ and we note that $r_{*}=-\infty$ is the black hole’s horizon. In hyperboloidal coordinates $(\tau,\rho)$ the domain is $\rho\in[\rho_{L},s]$ where $r_{*}(\rho=s)=\infty$.
@@ C.2 $a=0$ with hyperboloidal layers
We first consider a special case: when $a=0$ (scalar waves on Schwarzschild) the matrix $E$ is positive definite for all $\rho\in D\setminus\{s\}$. Upon further setting $M=0$, the results of this subsection also apply to the ordinary flatspace wave equation in hyperbolodial coordinates (44).
@@ C.3 The general case
We first show the diagonal entries are positive, which in turn requires that the diagonal entries of both $E_{\phi\phi}$ and $E_{\pi\pi}$ are positive. Because $E_{\phi\phi}=\mathbf{I}$ this is obviously true. Next we consider the diagonal of $E_{\pi\pi}$, whose elements are of the form $1-fc^{\ell}_{\ell m}$ where $f(r;a)$ is defined in Eq. (17). Let us first remark that $\Delta>0$ (and hence $f>0$) outside of the horizon, which our computational grid is restricted to. From App. A recall that $10<c^{\ell}_{\ell m}<$ 1, and so $1-f<1-fc^{\ell}_{\ell m}$. At this point, we use Mathematica to find an upper bound on $f(r;a)$. We do this by first finding $r=r_{\mathrm{max}}$ that maximizes $f(r;a)$ for a fixed value of $a$. Setting $\partial_{r}f=0$, we find $r_{\mathrm{max}}$ solves the cubic equation $r^{3}-3Mr^{2}+a^{2}r+a^{2}M=0$, which has one real root outside of the horizon. We now view $f(r=r_{\mathrm{max}};a)$ as a function of $a$. Numerical evidence (and physical intuition) suggests that this function is maximized at $a=1$. In fact, for $a=M=1$ we have $r_{\mathrm{max}}=1+\sqrt{2}$. Evaluating we get
@@ C.4 System at infinity
And so the norms of $C^{-1}\hat{A}C^{-1}$ and $A$ are identical at every point $\rho\in D$. In particular, at $\rho=s$ we have that $C^{-1}\hat{A}C^{-1}$ is bounded because $A$ is bounded. Therefore $C^{-1}\hat{A}C^{-1}$ exists, and hence the associated system (61) is symmetric hyperbolic everywhere in the the physical domain of interest, including $\rho=s$.
As shown in Fig. 2, our preferred system (which arises from using hyperboloidal layers with the reduction variables $\pi_{\ell m}=-\partial_{t}\psi_{\ell m}$ and $\phi_{\ell m}=\partial_{r*}\psi_{\ell m}$) exhibits the spectral convergence to round-off. Meanwhile, if we had instead used $\pi_{\ell m}=-\partial_{\tau}\psi_{\ell m}$ and $\phi_{\ell m}=\partial_{\rho}\psi_{\ell m}$ as our first-order reduction variables, does not show the expected spectral convergence. The next appendix considers this system in more detail.
@@ Appendix D An alternative first-order reduction
Throughout the main body of this paper, we have worked with a first-order reduction of Eq. (II.1) that arise from the reduction variables $\pi_{\ell m}=-\frac{\partial\psi_{\ell m}}{\partial t}$ and $\phi_{\ell m}=\frac{\partial\psi_{\ell m}}{\partial r_{*}}$ followed by a coordinate transformation 18. Appendix C shows the conditions under which the resulting system is positive (semi-definite) symmetric hyperbolic.
@@
where in carrying out the first-order reduction, we replaced the mixed-partial derivative term as $2\dot{\psi}^{\prime}_{\ell m}=-\partial_{\rho}\pi_{\ell m}+\partial_{\tau}\phi_{\ell m}$ and did not consider alternative choices. For example, the choice $2\dot{\psi}^{\prime}_{\ell m}=-2\partial_{\rho}\pi_{\ell m}$ will lead to a simpler $A$ matrix (cf. 25) but was not considered in our first-order reduction. If $\psi_{\ell m}$ solves the first-order system (71) it also solves the original second-order equation (D) provided $\phi_{\ell m}=\partial_{\rho}\psi_{\ell m}$.
Let us now put Eq. (71) into matrix form (25). For concreteness, consider the even $\ell$-mode sector with $m=0$ and $\ell_{\rm max}=2$. Then we have $U=\left[\pi_{00},\pi_{20},\phi_{00},\phi_{20}\right]^{T}$, and
@@
D.1 Potential issues with a singular boundary matrix
Hyperbolic initial-boundary problems (IBVPs) are typically called well-posed if a (weighted) $L_{2}$ norm of $U$, for example the energy ${\cal E}(\tau)$ define in Eq. (62), can be appropriately bounded by the initial data and boundary conditions Sarbach and Tiglio (2012); Majda and Osher (1975); Gustafsson et al. (1995). The IBVP is called strongly well-posed if one can also control the norm of the solution (and its regularity) in a neighborhood of the boundary; see Eq. (179) of Ref. Sarbach and Tiglio (2012).
Following the notation of Ref. Sarbach and Tiglio (2012), and noting that our system has one spatial dimension, the right boundary matrix is given by $P=A(\rho=s)$. Notice that some of the eigenvalues of the boundary matrix are zero. Hence, the boundary matrix is singular and, as a result, we cannot immediately conclude strong well-posedness even for symmetric hyperbolic systems Sarbach and Tiglio (2012).
Numerically, we have been unable to achieve good results with the system arising from reduction variables (70). For example, in the very simple case of $M=0$ (ie flatspace) the numerical solution to this system stops converging after about 4 to 5 digits of accuracy. This is shown most clearly in Fig. 2: compare “System 2” (the problematic one using reduction variables (70)) and “System 1” (the first-order system that arises from the reduction variables (15)). A singular boundary matrix could explain these empirical observations and further analysis following Refs. Ohkubo (1981); Rauch (1985); Secchi (1995) could be considered in future work. However, as mentioned above and repeated here, we caution the reader that because we do not understand the observed problematic behavior, we cannot rule out more pedestrian explanations like a code bug despite carefully checking our code.